OpenStudy (anonymous):

A tuition scale for a state university shows an increase of $1,300 each year. If the initial cost was $21,000, what will the cost of tuition be in twenty three years? In complete sentences, explain how an arithmetic sequence was applied

Parth (parthkohli):

Oh this one is easy :)

Parth (parthkohli):

This is an arithmetic sequence which goes like 21,000 ; 22,300 ; 23,600 and so on.

Parth (parthkohli):

If you notice, this is more like an arithmetic sequence with \(a_1\) as 21,000 and difference as \(1,300\).

Parth (parthkohli):

Using the formula \(a_n = a_1 + ( n - 1)d\), we can say that the money after 23 years would be \(a_{23} = 21000 + 1300( 23 - 1)\)

OpenStudy (anonymous):

ok but how wuld i solve it from there?

Parth (parthkohli):

\( \color{Black}{\Rightarrow a_{23} = 21000 + 1300(22) }\) \( \color{Black}{\Rightarrow a_{23} = 21000 + 28600 }\) Add them

OpenStudy (anonymous):

i got 49600

Parth (parthkohli):

Correct :D